ReaganLily:

The following dot plots represent the scores on the Chapter 5 quiz for Mrs. Chin's 4th and 6th period classes. Calculate the mean and mean absolute deviation (rounded to the nearest tenth) for both classes. Use your answer calculations from part A to answer all of the following questions: Which class period, on average, scored better on the quiz? By how much did they score better? How does the difference between the mean scores compare to the mean absolute deviation? Is there much overlap in the data? Write your answers in complete sentences.

ReaganLily:

1 attachment
Cupcake123456:

do you know witch number is the mean

ReaganLily:

no, I suck at math. I'm pretty bad when it comes to things like this

ReaganLily:

any chance someone could help me alittle?

Cupcake123456:

the mean is the average of the numbers

Cupcake123456:

add up all the number then divide

Cupcake123456:

many number there are

ReaganLily:

so to start,4th period.

ReaganLily:

divided by the amount of points?

Cupcake123456:

1st your gonna put the point in order from least to greatest

Cupcake123456:

once you do that check and see if theres an even number in the data set ..

ReaganLily:

what do you mean by put the point in order from least to greatest. and what graph the 4th period or 6th or what one are we working on right now Sorry if i'm asking alot of questions I'm really bad at math

Cupcake123456:

chapter 5 math quiz 4th period ,of course .

ReaganLily:

ok so I just list out the numbers in least to greatest?

Cupcake123456:

Yes , and after you do that you check for any even numbers.

ReaganLily:

like this? 10, 12, 12, 12, 13, 13, 13, 13, 14, 14, 15, 15, 15, 15, 15, 15, 16, 16

Cupcake123456:

where did those numbers come from ?

ReaganLily:

the picture above..

ReaganLily:

the plot that I attached

dude:

Yeah you're missing one number

ReaganLily:

what number? I'm actually so confused, sorry

Cupcake123456:

18

dude:

1551295984-5c76e5f024fa11494105de00-image.png

dude:

You missed 19

ReaganLily:

wait I'm confused I thought I was just supposed to do that for the top part (4th period, I thought I was only getting numbers for the first. )

ReaganLily:

oh right ugh one second.

dude:

Yeah this is the top part

ReaganLily:

10, 12, 12, 12, 13, 13, 13, 13, 14, 14, 15, 15, 15, 15, 15, 15, 16, 16, 19

Cupcake123456:

yeah my bad i meant 19 & that is the top part

dude:

Good now add them up

ReaganLily:

267?

Cupcake123456:

wait 1st are there any even numbers ?

ReaganLily:

wdym?

Cupcake123456:

What i mean is are there any even numbers ? do you know what even numbers are ?

ReaganLily:

I know what even numbers are just confused. like 10, 12 or like the number of dots? I'm just not all that good at all of this sorry I feel incredibly dumb

Cupcake123456:

ooopps my bad i was doing median xd

dude:

Actually count the number of dots there are

ReaganLily:

19

dude:

Ye Add them and then divide the answer by 19

Cupcake123456:

@dude thank you thats what i was tryna say

ReaganLily:

14

ReaganLily:

or 14.05 and so on

Cupcake123456:

once you add it it should come up to 267 , then u divide 267 by 19

Cupcake123456:

and yes your correct u should get 14.05

ReaganLily:

ok

ReaganLily:

that's the mean?

Cupcake123456:

14.05

ReaganLily:

rounded to the nearest tenth or is that just for the MAD

Cupcake123456:

not sure @dude

ReaganLily:

I'm guessing it means round both

dude:

Yeah

ReaganLily:

ok so then it's 14.1 is the mean?

ReaganLily:

and if that is true then I also figured out the mean of the other aswell. now for the MAD

ReaganLily:

now I don't remember nearly at all what the MAD is or how to find it

dude:

Not sure if they taught you with a certain notation MAD = \(\large \frac{\sum |xi-\overline{x}|}{n}\)

dude:

Do you know what each variable represents?

dude:

Or does this seem new to you?

ReaganLily:

completely new.

dude:

xi is a number in the set x with the line is the mean n is the total number of numbers {number of data}

dude:

The funny shaped E is just sum

ReaganLily:

huh...

ReaganLily:

this is for figuring out the MAD?

dude:

Yeah Do you have notes? (It would help since I was not taught the same way as you were)

ReaganLily:

no I don't.

ReaganLily:

I was going back through the lessons but I have no idea.

dude:

10, 12, 12, 12, 13, 13, 13, 13, 14, 14, 15, 15, 15, 15, 15, 15, 16, 16, 19 Subtract each point from the mean So... 10-14.1 12-14.1 12-14.1 and so on

ReaganLily:

ok

dude:

The || means to keep the number as a positive number So |1-5| = 4, not -4

ReaganLily:

do I have to repeat for the amount of points? so for say 12 how there is 2 points repeat it twice? this is going to get long....

dude:

Yes repeat for every point, even if it is the same point

ReaganLily:

10 - 14.1 = -4.1 12 - 14.1 = -2.1 12 - 14.1 = -2.1 12 - 14.1 = -2.1 13 - 14.1 = -1.1 13 - 14.1 = -1.1 13 - 14.1 = -1.1 13 - 14.1 = -1.1 14 - 14.1 = -0.1 14 - 14.1 = -0.1 15 - 14.1 = 0.9 15 - 14.1 = 0.9 15 - 14.1 = 0.9 15 - 14.1 = 0.9 15 - 14.1 = 0.9 15 - 14.1 = 0.9 16 - 14.1 = 1.9 16 - 14.1 = 1.9 19 - 14.1 = 4.9 hope I did this right...

dude:

Make all the negatives positives but yes

ReaganLily:

oh so instead of like -4.1 I do just 4.1

dude:

10 - 14.1 = +4.1 12 - 14.1 = +2.1 12 - 14.1 = +2.1 12 - 14.1 = +2.1 13 - 14.1 = +1.1 13 - 14.1 = +1.1 13 - 14.1 = +1.1 13 - 14.1 = +1.1 14 - 14.1 = +0.1 14 - 14.1 = +0.1 15 - 14.1 = 0.9 15 - 14.1 = 0.9 15 - 14.1 = 0.9 15 - 14.1 = 0.9 15 - 14.1 = 0.9 15 - 14.1 = 0.9 16 - 14.1 = 1.9 16 - 14.1 = 1.9 19 - 14.1 = 4.9

dude:

Ye

ReaganLily:

oh ok

ReaganLily:

so now?

dude:

Now all all the answers you got

dude:

add

ReaganLily:

ok

ReaganLily:

I think 29.1

ReaganLily:

I hope...

dude:

Yep and now divide by the number of dots

ReaganLily:

which was 19 right?

dude:

Ya

ReaganLily:

1. 53 and so on

dude:

Round to the nearest tenth

ReaganLily:

1.5

ReaganLily:

I just realized I have to do this all over again for the other period... omg

dude:

Yeah

dude:

But you get the idea now, right?

ReaganLily:

yes. for the most part. and then just the questions at the end.

ReaganLily:

is all that's left once I figure out the other MAD

dude:

Find the mean and MAD for the second set okay? First set: Mean = 14.1 MAD = 1.5

ReaganLily:

ok

ReaganLily:

and for the other now.

ReaganLily:

13 14 15 15 15 16 16 17 17 17 17 18 18 18 18 18 19 20 so then I do what again?

dude:

Add them up and divide by the number of dots in that set

ReaganLily:

6.7 is what I have for the mean fo the 2nd part

ReaganLily:

i mean no 16.7

ReaganLily:

and it's the mean that subtract everything from right?

ReaganLily:

13 - 16.7 = +3.7 14 - 16.7 = +2.7 15 - 16.7 = +1.7 15 - 16.7 = +1.7 15 - 16.7 = +1.7 16 - 16.7 = +0.7 16 - 16.7 = +0.7 17 - 16.7 = 0.3 17 - 16.7 = 0.3 17 - 16.7 = 0.3 17 - 16.7 = 0.3 18 - 16.7 = 1.3 18 - 16.7 = 1.3 18 - 16.7 = 1.3 18 - 16.7 = 1.3 18 - 16.7 = 1.3 19 - 16.7 = 2.3 20 - 16.7 = 3.3 hope this is right too.

ReaganLily:

and then added up is 26.2

ReaganLily:

and then I divide by the amount of dots, which is 18 in this case right?

dude:

Good

ReaganLily:

which then goes to 1.45 which rounded to the nearest tenth makes it 1.5

ReaganLily:

so 1.5 is the MAD for this one too?

dude:

Yep

ReaganLily:

ok so now on to the questions at the end?

ReaganLily:

Which class period, on average, scored better on the quiz? By how much did they score better? How does the difference between the mean scores compare to the mean absolute deviation? Is there much overlap in the data? these are them...

ReaganLily:

so for the first question. the 6th period is better right?

ReaganLily:

is it decided by the mean

dude:

Average is another word for mean (Which mean was greater?) (By how much?) (How are the means different to their MAD’s?)

ReaganLily:

What I have so far. 4th period: 14.1 is the mean, the MAD is: 1.5 6th period: 16.7 is the mean, the MAD is: 1.5 Which class period, on average, scored better on the quiz? Answer: 6th period got better scores. By how much did they score better? Answer: 2.6 Then the rest is to be done.

Cupcake123456:

what u need help with ?

ReaganLily:

just figuring out these last 2.

Cupcake123456:

witch 2 ?

ReaganLily:

the 2 above I said I haven't done yet.

Cupcake123456:

4th period: 14.1 is the mean, the MAD is: 1.5 6th period: 16.7 is the mean, the MAD is: 1.5 Which class period, on average, scored better on the quiz? Answer: 6th period got better scores. By how much did they score better? Answer: 2.6 Then the rest is to be done.

Cupcake123456:

this

ReaganLily:

that is what I have done. yes the rest is what I have

Cupcake123456:

im on my own math rn

ReaganLily:

ok so to start this one. How does the difference between the mean scores compare to the mean absolute deviation? this is something I'm confused about and don't know how or what the answer is.

ReaganLily:

What I have so far. 4th period: 14.1 is the mean, the MAD is: 1.5 6th period: 16.7 is the mean, the MAD is: 1.5 Which class period, on average, scored better on the quiz? Answer: 6th period got better scores. By how much did they score better? Answer: 2.6 the last question which is one above is tied into everything btw

ReaganLily:

if anyone knows anything about the last question above the what I have so far please let me know